# Physical partisan proximity outweighs online ties in predicting US voting outcomes

### v1

@@ Abstract
Affective polarization and increasing social divisions affect social mixing and the spread of information across online and physical spaces, reinforcing social and electoral cleavages and influencing political outcomes. Here, using aggregated and de-identified co-location and online network data, we investigate the relationship between partisan exposure and voting patterns in the USA by comparing three dimensions of partisan exposure: physical proximity and exposure to the same social contexts, online social ties, and residential sorting. By leveraging various statistical modeling approaches, we consistently find that partisan exposure in the physical space, as captured by co-location patterns, more accurately predicts electoral outcomes in US counties, outperforming online and residential exposures across metropolitan and non-metro areas. Moreover, our results show that physical partisan proximity is the best predictor of voting patterns in swing counties, where the election results are most uncertain. We also estimate county-level experienced partisan segregation and examine its relationship with individuals’ demographic and socioeconomic characteristics. Focusing on metropolitan areas, our results confirm the presence of extensive partisan segregation in the US and show that offline partisan isolation, both considering physical encounters or residential sorting, is higher than online segregation and is primarily associated with educational attainment. Our findings emphasize the importance of physical space in understanding the relationship between social networks and political behavior, in contrast to the intense scrutiny focused on online social networks and elections.
keywords:
political polarization, social networks, computational social science
@@ Introduction
Affective polarization has emerged as a critical concern for contemporary democracies, dividing populations along partisan and political lines and impacting various aspects of individuals’ lives [1, 2, 3, 4]. Moreover, recent years have witnessed considerable attention on the impact of online social media and news consumption - specifically misinformation and disinformation - on opinion dynamics, political polarization and, consequently, on political elections [5, 6, 7, 8, 9, 10, 11, 12, 13]. However, political polarization and partisan exposure cannot be solely explained by human behavior on social media platforms, as they are influenced by various dimensions of individuals’ daily lives, including physical and online environments, as well as the role of social networks and existing social divisions.
@@
While social mixing and diversity are associated with positive effects on many socioeconomic aspects of human life [14, 15, 16], social segregation in physical spaces has been observed concerning ethnicity and migrants’ characteristics [17], income inequality [18, 19], and partisan affiliation [20, 21, 22, 23]. Such divisions might impact the diversity of individuals’ social exposure and the spread of information, contributing to the maintenance of social and electoral cleavages. While previous studies have primarily focused on partisan exposure in terms of residential locations, experienced social exposure extends beyond the place of residence and involves routines, habits, and social interactions throughout people’s daily experiences [24, 19, 25, 26], both in physical and online spaces [27]. In this context, social networks represent and shape many aspects of individuals’ lives, with relevant impacts on the access to novel information [28, 29, 30], economic actions and prosperity [31, 14, 32]. Positive associations have also been observed between geographic areas, where regions with higher levels of connection generally have better social and economic outcomes [33, 34]. From this perspective, understanding partisan social exposure and segregation dynamics is crucial for studying the transmission of political choices through social networks, following social influence and voting contagion dynamics [35, 36]. However, a comparative analysis between online and physical partisan exposure in social networks, and their relationship with political voting patterns is still lacking.
@@
In this context, our study aims to understand the relationship between social exposure and voting outcomes by comparing partisan exposure in online and physical spaces. To this aim, we explore the relationship between partisan exposure and voting patterns in US counties by estimating the association between political vote and physical proximity and exposure to the same social context, online social connections, and residential sorting. Specifically, first, we ask whether there are differences between partisan exposure across the different dimensions, both in metropolitan and non-metropolitan areas. Second, we evaluate the dimension that best predicts voting patterns across the urban-rural divide and historical voting outcomes. Third, we investigate how partisan segregation relates to demographic and socioeconomic characteristics of US counties.
@@
To compare the online and physical exposure patterns, we leverage three datasets. First, we consider offline social connectedness using the Colocation Maps dataset, which provides the co-location probability between two randomly selected individuals residing in different US counties [37]. Second, we analyze the Social Connectedness Index (SCI) dataset, which measures the relative probability of friendship on Facebook between two randomly selected individuals residing in different counties [33]. Third, we consider the probability of being exposed to Democratic and Republican voters based on the place of residence, by looking at individuals’ neighbors, using a publicly available dataset [38]. We compute partisan exposure to Democrats and Republicans for each county in the contiguous United States and each dimension: offline, online, and residential. We then model the relationship with voting patterns at the county level and seek to determine which dimension best explains the observed variance of voting patterns.
@@
Our findings reveal the strong and significant role of physical proximity and local exposure in predicting the voting outcomes of the US counties, in both metropolitan and non-metro areas. Moreover, when focusing on the electoral outcomes of swing counties, where political results are less predictable from year to year, we show that physical proximity significantly outweighs all other dimensions in the prediction task. Finally, we find that partisan segregation is higher in offline social networks than online ones, with physical partisan segregation primarily associated with the county population’s educational attainment and the urban-rural divide. Overall, our work contributes to a better understanding of the role of social exposure in determining political outcomes in the US, specifically the voting patterns of the US presidential elections at the county level. Our results highlight the importance of gauging the relative role of different types of exposure, offline and online, when considering their effects on election results.
@@ Results
Leveraging three datasets representing different dimensions of individuals’ daily lives, we compare partisan exposure and its relationship with voting behavior at the county level in the United States. We first compute partisan exposure from the Colocation Maps[37], the Social Connectedness Index[33], and at the residential level[23]. We then estimate the relationship between partisan exposure across these dimensions and the voting patterns of US counties. Finally, we investigate how partisan segregation relates to individuals’ demographic and socioeconomic characteristics in metropolitan areas.
Partisan exposure across different dimensions
We first compute the relative exposure to Republican and Democratic voters for each dimension. Physical partisan exposure is derived from the Colocation Maps, while online exposure is measured from the Social Connectedness Index (SCI). Specifically, the former provides the co-location (Fig. 1a) probability between two randomly selected individuals from US counties $i$ and $j$ [37]. A co-location event (Fig. 1a) is registered when two Facebook users are in the same location for at least 5 minutes [37]. Conversely, the SCI dataset provides the relative probability of friendship on Facebook between two users from $i$ and $j$ (Fig. 1a) [33]. We refer to the Materials and Methods Section for further details.
Despite a high correlation between Colocation Maps and SCI raw probabilities (Pearson’s $r=0.83$), their corresponding networks show significant structural differences. Specifically, as shown in Figs. 1b and 1c, which compare the connection probabilities for Jackson County, MO, online connections are more homogeneously distributed across the country than offline co-location events. This is further explored in the SI Appendix, Fig. S3, which shows that US counties display a higher heterogeneity of social connections (network diversity $D(i)$) in the online space, with a median value of $0.39$ compared to $D(i)=0.15$ offline. Moreover, US counties have higher online extroversion (defined as the ratio between external and internal connection probabilities) than offline extroversion, with a median value of $1.19$ compared to $0.27$.
Experienced physical and online partisan exposures, $PE$, for a county $i$ are computed as the relative exposure to Republican and Democratic voters of a county $j$ weighted by either the co-location or online connection probabilities between $i$ and $j$ (see Materials and Methods Section for more details). Specifically, we define the share of Republican and Democratic voters as the average share of votes for Republican and Democratic candidates in the last three presidential elections (i.e., 2012, 2016, and 2020), following the normal vote concept [39] to account for candidate- or election-specific influences on voter turnout and election results. Partisan exposure ($PE$) is a measure that ranges from 0 (low exposure to either Democratic or Republican voters) to 1 (high exposure to either Democratic or Republican voters).
@@
We then compute residential partisan exposure for each county from the dataset provided by R. Enos [38, 23], which provides the conditional probabilities of being exposed to either Republican or Democratic voters for both Republicans and Democrats at the residential level, based on the proximity to the nearest 1,000 individuals who have registered to vote. To the aim of our study, we compute partisan exposure to Democratic and Republican voters for each US county, as the probability that a random individual is exposed to Republicans or Democrats. Further details can be found in the Materials and Methods Section.
[…]
@@
Partisan exposure and voting behavior
@@
We estimate the relationship between physical, online, and residential partisan exposure and voting behavior through several statistical models to assess their relative contribution to voting dynamics. As with the computation of partisan exposure, voting patterns of US counties are defined as the average share of Democratic and Republican voters in the last three presidential elections (i.e., 2012, 2016, and 2020). The shares are not complementary because we account for third-party and null votes. According to the normal vote concept [39], this method adjusts for candidate- or election-specific influences on voter turnout and election results, providing a more accurate assessment of voting patterns. In all the models, the shares of Democrats and Republicans in the US counties are our dependent or target variables.
@@
We first model each dimension of partisan exposure ($PE$) and voting patterns for Democrats and Republicans separately with spatial autoregressive lag models [40], accounting for spatial autocorrelation across all counties in the contiguous United States. We use k-nearest neighbor to compute spatial weights, with different values of $k$ ($5$, $7$, and $10$) for robustness. Similarly, we evaluate the relationship between partisan exposure and voting patterns in metropolitan and non-metropolitan areas with ordinary least squares (OLS) regressions. Finally, we compare swing and non-swing counties, considering counties with varying political outcomes in the last three presidential elections and whose voting patterns are consequently less predictable.
@@
In Fig. 2, we show the results for both spatial and OLS models, for Democrats and Republicans, further disaggregated by metropolitan and non-metropolitan counties. We compare the models’ ability to explain the variance in the dependent variable using $R^{2}$, with consistent results when measuring the model’s quality with the Akaike Information Criterion (AIC) and log-likelihood. Considering all US counties and accounting for spatial autocorrelation, physical partisan exposure to both Democrats and Republicans, as captured by Colocation Maps, is the dimension that best explains the variance in the share of Democratic ($R^{2}=0.97$) and Republican ($R^{2}=0.97$) votes respectively, when compared to online ($R^{2}=0.87$ and $0.85$) and residential ($R^{2}=0.80$ and $0.75$) exposures (Fig. 2a). Moreover, when comparing online and residential partisan exposures, the former has a higher explanatory power of the variance in voting patterns than the latter, for Democrats and Republicans. Disentangling the analysis across the urban-rural axis according to the RUCC, we find that while in non-metro areas offline and online partisan exposures have similar predictive power ($R^{2}=0.98$ and $0.94$ respectively for Republicans and $R^{2}=0.99$ and $0.95$ for Democrats), in metro areas the partisan exposure captured by Colocation Maps significantly outperforms both the other dimensions, with the $R^{2}$ of the model related to offline exposure equal to $0.93$ (Republicans) and $0.92$ (Democrats) compared to online ($0.73$ and $0.76$ respectively) and residential ($0.80$ and $0.69$) exposures (Figs. 2b and 2c). Therefore, in metropolitan areas, voting patterns are predominantly associated with the partisan exposure to Democrats and Republicans in the physical space, when sharing the same social context.
@@
To check the robustness of our results, we test our findings with different modeling approaches. First, we perform a dominance analysis [41] to compare the three dimensions, accounting for possible demographic and socioeconomic confounders, including the shares of Hispanics and Latinos, African Americans, graduated individuals, unemployment, and urban population. These characteristics are selected based on the literature on the determinants of political outcomes [42, 43, 44, 42, 45, 46] and by computing Variance Inflation Factors (VIFs) to address multicollinearity issues. We refer to the Materials and Methods Section for further details about the variable selection. From this perspective, dominance analysis allows us to deal with highly correlated variables and determine the relative importance of each predictor. Fig. 2d confirms the results obtained using the spatial lag models (Fig. 2a), with the higher relative importance of offline exposure in explaining the variance of US voting patterns. Specifically, offline exposure shows greater relative importance for both Republicans ($35.29\%$) and Democrats ($36.12\%$) compared to the online ($27.43\%$ and $29.5\%$ respectively) and residential ($20.11\%$ and $15.97\%$) exposures. Moreover, the results of the OLS models (Figs. 2b and 2c) are confirmed by the dominance analysis in both metropolitan (Fig. 2e) and non-metropolitan areas (Fig. 2f). See the SI Appendix for complete results.
@@
Second, we enhance the generalizability of our findings by modeling the relationship between the three $PE$ dimensions and the voting patterns using the Elastic Net models with k-fold cross-validation [47] (see Materials and Methods Section and SI Appendix, Table S22). Consistently with previous results, physical exposure, as captured by Colocation Maps, has the greatest impact on the prediction compared to online and residential exposures ($\beta=1.08$ compared to $-0.18$ and $0.09$ for Republicans and $\beta=1.15$ compared to $-0.22$ and $0.06$ for Democrats). Overall, the models achieve high performances in terms of generalizability, with $R^{2}$ equal to $0.97$ in the test sets for both models.
Figure 3: Relative contribution of the three dimensions of partisan exposure on voting patterns comparing swing and non-swing counties. (a and b) $R^{2}$ of the OLS models for the swing counties and non-swing counties, respectively, highlight the higher predictive power of physical exposure, especially in the swing counties. (c and d) Dominance analysis for the swing and non-swing counties, respectively, considering the three dimensions, along with demographic and socioeconomic controls. (e and f) Map of the swing counties in the contiguous US and distribution according to the RUCC classification.
We then categorize counties into swing and non-swing counties based on their voting patterns in the presidential elections from 2012 to 2020. We also consider the 2008 presidential election as an additional robustness check in the SI Appendix (Fig. S2). Swing counties are defined as counties that experienced electoral volatility, shifting their majority support among different parties across three consecutive presidential elections. We refer to the Materials and Methods Section for further details. The experienced physical exposure $PE$ to Democrats and Republicans best explains the variance in voting patterns for both swing and non-swing counties. However, when comparing swing and non-swing counties, the difference between the predictive power of physical and online exposures is larger in the swing counties in both OLS models and dominance analysis. Specifically, in the OLS models for Republicans and Democrats, physical exposure in swing counties explains $58\%$ and $57\%$ of the variance, respectively, while online exposure accounts for only $16\%$ and $21\%$ of the variance, respectively (Fig. 3a and SI Appendix, Tables S10 and S18). In contrast, in non-swing counties, the differences between offline and online predictive power are reduced to $11\%$ in the model for the Republican vote and $9\%$ in the model for the Democratic vote (Fig. 3c and SI Appendix, Tables S11 and S19). Therefore, voting patterns are strongly associated with partisan exposure in the physical space, demonstrating a greater influence than online connections and residential proximity.

### v2

@@ Abstract
Affective polarization and increasing social divisions affect social mixing and the spread of information across online and physical spaces, reinforcing social and electoral cleavages and influencing political outcomes. Here, using individual survey data and aggregated and de-identified co-location and online network data, we investigate the relationship between partisan exposure and vote choice in the US by comparing offline and online dimensions of partisan exposure. By leveraging various statistical modeling approaches, we consistently find that partisan exposure in the physical space, as captured by co-location patterns, more accurately predicts electoral outcomes in US counties, outperforming online and residential exposures. Similarly, offline ties at the individual level better predict vote choice compared to online connections. We also estimate county-level experienced partisan segregation and examine its relationship with individuals’ demographic and socioeconomic characteristics. Focusing on metropolitan areas, our results confirm the presence of extensive partisan segregation in the US and show that offline partisan isolation, both considering physical encounters or residential sorting, is higher than online segregation and is primarily associated with educational attainment. Our findings emphasize the importance of physical space in understanding the relationship between social networks and political behavior, in contrast to the intense scrutiny focused on online social networks and elections.
@@ Introduction
Affective polarization has emerged as a critical concern for contemporary democracies, dividing populations along partisan and political lines and impacting various aspects of individuals’ lives [1, 2, 3, 4, 5]. Moreover, recent years have witnessed considerable attention to the impact of online social media and news consumption - specifically misinformation and disinformation - on opinion dynamics, political polarization, and, consequently, on political elections [6, 7, 8, 9, 10, 11, 12, 13, 14]. However, political polarization and partisan exposure cannot be solely explained by human behavior on social media platforms, as they are influenced by various dimensions of individuals’ daily lives, including physical and online environments, the role of social networks, and existing social divisions.
Homogeneous social networks can contribute to increased affective polarization and the maintenance of social and electoral cleavages through interactions with like-minded individuals [15, 3]. In physical spaces, the debate has focused on the division of partisans at the residential level, where the grouping of like-minded people has increased at the geographical level in the United States [16, 17, 18]. Specifically, individuals generally experience limited exposure to outgroup partisans, even within the same city or neighborhood [17]. Furthermore, geographical sorting is significant both within and between geographical units [18], with larger differences observed within areas than between them [19]. However, in terms of residential mobility, although Americans prefer politically compatible communities [20], they do not relocate for political reasons [21]. Additionally, when considering individual mobility behavior, experienced partisan segregation in activity spaces varies depending on the physical space and geographic area [22], and online connections between geographic areas are driven by partisan homophily [23]. Finally, individual survey data shows that political polarization has increased in association with offline social network homophily [24]. On the other hand, in online spaces and social media, affective polarization may be associated with self-selection and echo chambers [7, 25] or with the exposure to nonlocal interactions that may fuel political conflicts [12]. However, ideological segregation in the US is higher in face-to-face interactions than in offline and online news consumption [26], and partisan segregation in news consumption appears to be higher in traditional media than in online social networks [26, 13], with greater heterogeneity in partisan exposure online [27]. Moreover, polarization has increased most significantly among social and demographic groups that are less likely to use social media [10], while deactivation experiments have shown no significant effects on voter turnout in US presidential elections or affective and ideological polarization [14].
@@
Although previous studies have focused mainly on partisan exposure in terms of residential locations, experienced social exposure extends beyond the place of residence and involves routines, habits, and social interactions throughout people’s daily experiences [28, 29, 22, 30, 31, 32], both in physical and online spaces [33]. In this context, social networks represent and shape many aspects of individual lives, with relevant impacts on access to novel information [34, 35, 36], economic actions and prosperity [37, 38, 39], and political participation [40, 41, 42]. In particular, the literature highlights how social networks, including spouses, family, close friends and acquaintances, and the social environment foster political discussion and participation [43, 44, 45, 46, 47], and reinforce political beliefs [48, 40, 49]. Moreover, some studies have shown that individuals’ vote choices are associated with the voting preferences of their political discussion networks [49, 50]. Social exposure and its influence on political behavior is represented by both context and contact. Intergroup contact varies in depth, from casual to cooperative or selective encounters, and duration, from brief to sustained exposure [51]. While most studies focus on cooperative and sustained interactions and draw on contact theory [52], research shows that proximity and casual exposure can also have significant effects on political behavior [53, 54]. Understanding partisan social exposure and the dynamics of segregation in both physical and online dimensions is crucial to studying the transmission of political choices through social networks, following social influence and voting contagion [55, 56]. However, a large-scale comparative analysis between online and physical partisan exposure in social networks and their relationship with political behavior is still lacking.
@@
In this context, our study aims to understand the relationship between partisan exposure in online and physical spaces and voting outcomes. To this end, we integrate large-scale aggregated network data and individual surveys. Specifically, we first investigate whether partisan exposure differs across physical proximity and exposure to the same social context, online social connections, and residential sorting, both in metropolitan and non-metropolitan areas. Second, we explore how partisan segregation at the county level relates to the demographic and socioeconomic characteristics of US counties. Third, we evaluate the dimension that best predicts voting patterns at both the county and individual levels. Our study is not intended to predict the results of future presidential elections. Instead, it focuses on contrasting existing online social networks with those in the physical world to describe electoral patterns.
@@
To compare the online and physical exposure patterns, we leverage four datasets. First, we consider offline social connectedness using the Colocation Maps dataset, which provides the co-location probability between two randomly selected individuals residing in different US counties [57]. Second, we analyze the Social Connectedness Index (SCI), which measures the relative probability of friendship on Facebook between two randomly selected individuals who reside in different counties [58]. Third, we consider the probability of being exposed to Democratic and Republican voters based on the place of residence, looking at the neighbors of the individuals, using a publicly available dataset [59]. We compute partisan exposure to Democrats and Republicans for each county in the contiguous United States and each dimension: offline, online, and residential. We then model the relationship with voting patterns at the county level and seek to determine which dimension best explains the observed variance of voting patterns. Fourth, we used the 2020-2022 Social Media Study data [60] provided by the American National Election Studies (ANES) to model the relationship between partisan exposure to Democrats and Republicans and the political vote in the 2020 presidential elections at the individual level. These datasets capture different types of contact across distinct contexts (see SI, Table S1). Colocation Maps reflect physical co-presence, including both casual and selective, as well as brief and repeated encounters. These contacts are generally passive. The Social Connectedness Index captures digital social ties, typically selective, sustained, and active. Residential proximity represents the physical context in which all types of contact, casual or selective, brief or sustained, can occur. This form of contact is largely passive, shaped by where people live. Finally, survey data provide detailed measures of interpersonal relationships, reflecting selective, sustained, and primarily active forms of social contact.
@@
Our findings reveal the strong and significant role of physical proximity and local exposure in predicting the voting outcomes of the US counties, as well as the role of offline ties in predicting individuals’ political votes in the 2020 presidential election when compared to online ties. Moreover, we find that partisan segregation is higher in offline social networks than in online ones, with physical partisan segregation primarily associated with the county population’s educational attainment and the urban-rural divide. Overall, our work contributes to a deeper understanding of how social exposure reflects political outcomes in the US, specifically the voting patterns in presidential elections at both the county and individual levels, highlighting the effects of partisan segregation on voting polarization, with potential implications for affective polarization. Our results highlight the importance of gauging the relative role of different types of exposure, offline and online, when considering their effects on election results.
@@ Results
In this study, we compare offline and online partisan exposure and its relationship with vote choice at both the county and individual levels in the United States, and we examine partisan segregation. We first measure partisan exposure at the county level. Second, we analyze how partisan segregation relates to individuals’ demographic and socioeconomic characteristics within metropolitan areas. Finally, we estimate the relationship between partisan exposure and vote choice at both the county and individual levels.
Defining partisan exposure across different dimensions
We first compute the relative exposure to Republican and Democratic voters for each dimension. Physical partisan exposure is derived from the Colocation Maps, while online exposure is measured from the Social Connectedness Index (SCI). Specifically, the former provides the co-location probability (Fig. 1 a) between two randomly selected individuals from US counties $i$ and $j$ [57]. A co-location event (Fig. 1a) is registered when two Facebook users are at the same location for at least 5 minutes [57]. In contrast, the SCI dataset provides the relative probability of friendship on Facebook between two users from $i$ and $j$ (Fig. 1a) [58]. We refer to the Materials and Methods Section for further details.
Despite a high correlation between Colocation Maps and SCI raw probabilities (Pearson’s $\rho=0.83$), their corresponding networks show significant structural differences. Specifically, as shown in Figs. 1b and 1c, which compare the connection probabilities for Jackson County, MO, online connections are more homogeneously distributed across the country than offline co-location events. This is further explored in Supplementary Information, Fig. S1, which shows that US counties display a higher heterogeneity of social connections (network diversity $D(i)$) in the online space, with a median value of $0.39$ compared to $D(i)=0.15$ offline. Moreover, US counties have higher online extroversion (defined as the ratio between external and internal connection probabilities) than offline extroversion, with a median value of $1.19$ compared to $0.27$.
Experienced physical and online partisan exposures, $PE$, for a county $i$ are computed as the relative exposure to Republican and Democratic voters of a county $j$ weighted by either the co-location or online connection probabilities between $i$ and $j$ (see Materials and Methods Section for more details). Specifically, we define the share of Republican and Democratic voters as the average share of votes for Republican and Democratic candidates in three past presidential elections (that is, 2012, 2016 and 2020), following the normal vote concept [61] to account for candidate- or election-specific influences on voter turnout and election results. Partisan exposure ($PE$) is a measure that ranges from 0 (low exposure to Democratic or Republican voters) to 1 (high exposure to either Democratic or Republican voters).
@@
We then calculated the residential partisan exposure for each county from the dataset provided by R. Enos [59, 17], which provides the conditional probabilities of being exposed to Republican or Democratic voters for both Republicans and Democrats at the residential level, based on proximity to the nearest 1,000 individuals who have registered to vote. To the aim of our study, we compute the partisan exposure to Democratic and Republican voters for each US county as the probability that a random individual is exposed to Republicans or Democrats. Therefore, while measures of offline and online exposure are calculated from network ties between counties based on their political vote, residential exposure is derived from voter registration data. Further details can be found in the Materials and Methods Section.
To investigate the differences between partisan exposure across the three dimensions, we performed t-tests and Kolmogorov-Smirnov tests between the distributions of offline, online, and residential exposures (SI, Table S2). Note that the distributions are not population-weighted. Considering all counties in the contiguous United States (Fig. 1d), we find significant differences ($p<.001$) in partisan exposure to Republican voters when comparing residential exposure to online or offline exposure. However, no significant difference ($p>.05$) is observed between the means of the distributions of online and offline exposures (SI, Table S3). For Democratic voters, the differences in partisan exposure in the various dimensions are also significant ($p<.001$). Finally, we categorize counties into metropolitan and non-metro areas based on the Rural-Urban Continuum Codes (RUCC) and test the significance of the differences between partisan exposures (Fig. 1d; see SI, Tables S4, S5, and S6 for complete results).
Partisan segregation across demographic and socioeconomic factors
To investigate the levels of social segregation across political lines in the US, we estimate partisan segregation for each dimension as the net difference between exposure to Republicans and exposure to Democrats in a given county. The metric ranges from -1 (indicating exclusive exposure to Democrats) to 1 (indicating exclusive exposure to Republicans) and accounts for third-party voters. Intermediate values represent different mixing levels, with 0 indicating an equal share of exposure between the two electorates (see Materials and Methods Section for further details). We then investigated the relationship between partisan segregation and the demographic and socioeconomic characteristics of the county, focusing on metropolitan areas. To this end, we predict partisan segregation for each network dimension using Gradient Boosting (GB) regressions, based on a set of demographic and socioeconomic factors: the shares of Hispanics and Latinos, African Americans, graduated individuals, unemployment, and urban population. Furthermore, we improve the explainability of the model by computing SHAP values [62], which provide importance scores that represent the impact of each variable on the predicted outcome. We refer to Materials and Methods for further details on machine learning models.
Overall, partisan segregation in the physical world, both considering co-location maps and residential sorting, is higher than online segregation. As shown by the distributions of Fig. 2, in the offline and residential networks we observe counties characterized more frequently by $PS>0.5$ (highly Republican-segregated) or PS<0.5PS<-0.5 (highly Democrat-segregated). In particular, largely populated urban areas display values of PS<0.5PS<-0.5 in offline networks that are not met by online links.
Figure 2: Partisan segregation across demographic and socioeconomic factors for each dimension. Maps and population-weighted distributions of the partisan segregation as captured by the Colocation Maps (a), Social Connectedness Index (b), and at the residential level (c). The red solid line represents the weighted average, while the black dashed line indicates balanced social mixing (0). For each dimension, the SHAP values distributions, computed from the GB regressions, highlight how the demographic and socioeconomic characteristics of the counties impact partisan segregation. Predictors are ordered by their impact on the final prediction, with points colored red for high and green for low values.
In general, GB regressions effectively explain the variance of offline segregation ($R^{2}=0.63$), online ($0.55$), and residential ($0.64$). Moreover, Fig. 2 shows the distribution of SHAP values for each predictor related to offline, online, and residential partisan segregation in metropolitan areas. The predictors are ordered according to the impact on the prediction, with the most impactful variable at the top, and each point represents an observation. Finally, each point is colored according to its value, with red indicating high values and green indicating low values. The absolute impacts of the predictors are reported in the Supplementary Information (Table S7).
The percentage of graduated individuals is the best predictor for both partisan segregations computed from offline networks (average absolute impact equal to $0.096$) and online ($0.067$) networks. While counties with higher shares of graduated individuals tend to be exposed to Democrats (with higher Democratic segregation), counties with less educated individuals tend to be co-located and connected online with individuals who vote for the Republican party (Figs. 2a and 2b). In contrast, at the residential level, the educational level follows similar patterns but is the third predictor (mean average impact equal to $0.067$) in terms of variable importance (Fig. 2c).
Furthermore, partisan segregation is closely related to the presence of ethnic communities. Specifically, the presence of a large percentage of African Americans is the second most relevant predictor of partisan segregation both online ($0.037$) and at the residential level ($0.075$) in metropolitan areas. Large African American communities are associated with predominant exposure to Democrats, while areas with a low African American presence tend to Republican segregation (Fig. 2). Finally, as shown in Fig. 2, the presence of Latinos and Hispanics is the least relevant predictor on all dimensions, namely offline ($0.022$), online ($0.017$) and at the residential level ($0.029$). Interestingly, a larger share of Latinos and Hispanics in US counties are associated with higher levels of Democratic and Republican offline and online segregation.
@@
Offline partisan exposure better predicts the voting behavior of counties
@@
We estimate the relationship between physical, online and residential partisan exposure and vote choice through several statistical models to assess their relative contribution to voting dynamics. As with the calculation of partisan exposure, voting patterns of US counties are defined as the average share of Democratic and Republican voters in the last past presidential elections (i.e., 2012, 2016, and 2020). The shares are not complementary because we account for third-party and null votes. According to the normal vote concept [61], this method adjusts for candidate- or election-specific influences on voter turnout and election results, providing a more accurate assessment of voting patterns. In all the models, the shares of Democrats and Republicans in the US counties are our dependent or target variables.
@@
We first model each dimension of partisan exposure ($PE$) and voting patterns for Democrats and Republicans separately with spatial autoregressive lag models [63], accounting for spatial autocorrelation across all counties in the contiguous United States. We design distinct models due to the high correlation between the dimensions, as shown in the Supplementary Information (SI, Fig. S2). We use k-nearest neighbor to compute spatial weights, with different values of $k$ (55, 77, and $10$) for robustness. Similarly, we evaluate the relationship between partisan exposure and voting patterns in metropolitan and non-metropolitan areas with ordinary least squares (OLS) regressions. Finally, we measure direct, indirect, and total effects of the independent variables for spatial models (see SI, Table S11 and Fig. S3a), as well as marginal effects for metro and non-metro areas (Fig. S3b and S3c). See Supplementary Information for complete results (Tables S8 to S19).
@@
We estimate the relationship between physical, online and residential partisan exposure and vote choice through several statistical models to assess their relative contribution to voting dynamics. As with the calculation of partisan exposure, voting patterns of US counties are defined as the average share of Democratic and Republican voters in the last past presidential elections (i.e., 2012, 2016, and 2020). The shares are not complementary because we account for third-party and null votes. According to the normal vote concept [61], this method adjusts for candidate- or election-specific influences on voter turnout and election results, providing a more accurate assessment of voting patterns. In all the models, the shares of Democrats and Republicans in the US counties are our dependent or target variables.
We first model each dimension of partisan exposure ($PE$) and voting patterns for Democrats and Republicans separately with spatial autoregressive lag models [63], accounting for spatial autocorrelation across all counties in the contiguous United States. We design distinct models due to the high correlation between the dimensions, as shown in the Supplementary Information (SI, Fig. S2). We use k-nearest neighbor to compute spatial weights, with different values of $k$ (55, 77, and $10$) for robustness. Similarly, we evaluate the relationship between partisan exposure and voting patterns in metropolitan and non-metropolitan areas with ordinary least squares (OLS) regressions. Finally, we measure direct, indirect, and total effects of the independent variables for spatial models (see SI, Table S11 and Fig. S3a), as well as marginal effects for metro and non-metro areas (Fig. S3b and S3c). See Supplementary Information for complete results (Tables S8 to S19).
@@
In Fig. 3, we show the results for both spatial and OLS models, for Democrats and Republicans, further disaggregated by metropolitan and non-metropolitan counties. We compare the models’ ability to explain the variance in the dependent variable using $R^{2}$, with consistent results when measuring the model’s quality with the Akaike Information Criterion (AIC) and log-likelihood. Considering all US counties and accounting for spatial autocorrelation, physical partisan exposure to both Democrats and Republicans, as captured by Colocation Maps, is the dimension that best explains the variance in the share of Democratic ($R^{2}=0.97$) and Republican ($R^{2}=0.97$) votes, respectively, compared to online ($R^{2}=0.87$ and $0.85$) and residential ($R^{2}=0.80$ and $0.75$) exposures (Fig. 3a). Moreover, when comparing online and residential partisan exposures, the former has a higher explanatory power of the variance in voting patterns than the latter for Democrats and Republicans. Disentangling the analysis across the urban-rural axis according to the RUCC, we find that while in non-metro areas offline and online partisan exposures have similar predictive power ($R^{2}=0.98$ and $0.94$ respectively for Republicans and $R^{2}=0.99$ and $0.95$ for Democrats), in metro areas the partisan exposure captured by Colocation Maps significantly outperforms both the other dimensions, with the $R^{2}$ of the model related to offline exposure equal to $0.93$ (Republicans) and $0.92$ (Democrats) compared to online ($0.73$ and $0.76$ respectively) and residential ($0.80$ and $0.69$) exposures (Figs. 3b and 3c). Therefore, in metropolitan areas, voting patterns are predominantly associated with partisan exposure to Democrats and Republicans in the physical space, when sharing the same social context.
@@
To verify the robustness of our results, we test our findings with different modeling approaches. First, we perform a dominance analysis [64] to compare the three dimensions, taking into account possible demographic and socioeconomic confounders, including the shares of Hispanics and Latinos, African Americans, graduated individuals, unemployment, and urban population. These characteristics are selected based on the literature on the determinants of political outcomes [65, 66, 67, 65, 68, 69] and by computing the variance inflation factors (VIF) to address multicollinearity issues. We refer to the Materials and Methods Section for further details about the variable selection. From this perspective, dominance analysis allows us to deal with highly correlated variables (SI, Fig. S2) and determine the relative importance of each predictor. Fig. 3d confirms the results obtained using the spatial lag models (Fig. 3a), with the higher relative importance of offline exposure in explaining the variance of US voting patterns. Specifically, offline exposure shows greater relative importance for both Republicans ($35.29\%$) and Democrats ($36.12\%$) compared to online ($27.43\%$ and $29.5\%$, respectively) and residential ($20.11\%$ and $15.97\%$) exposures. Moreover, the results of the OLS models (Figs. 3b and 3c) are confirmed by the dominance analysis in both metropolitan (Fig. 3e) and non-metropolitan areas (Fig. 3f).
@@
Second, we improve the generalizability of our findings by modeling the relationship between the three dimensions of $PE$ and the voting patterns using Random Forest and Elastic Net models with k-fold cross-validation [70] (see Materials and Methods Section and SI, Tables S20 and S21). Consistently with previous results, physical exposure, as captured by Colocation Maps, has the greatest impact on prediction compared to online and residential exposures. The findings hold when considering both the variable importance in the Random Forest models and the coefficients of the Elastic Net models (see the SI, Section S7). In general, the models achieve high performance in terms of generalizability, with $R^{2}$ between $0.97$ and $0.98$ in the test sets for all models.
To compare our results with official statistics, we calculated partisan exposure using the 2016-2020 5-Year ACS Commuting Flows and modeled the relationship between partisan exposure and the voting patterns of the US counties. Partisan exposure, as captured by commuting flows, outperforms online and residential proximity in predicting voting patterns, but shows lower performance compared to physical exposure, as confirmed by both spatial models and dominance analysis (SI, Tables S22 to S25 and Fig. S4).